# Binary population synthesis of the Galactic canonical pulsar population

### v1

@@ 3.2 Results
Figure 2: CDFs of physical quantities produced by the COMPAS models described in Table. 1 (coloured), compared to the data from the catalogs (black). CDFs from a given model are drawn from a randomised sampling from 10 realisations. From top left to bottom right, the panels show the distributions of $\gamma$-ray pulsar $\dot{E}$, radio pulsar $\dot{E}$, $\gamma$-ray flux, radio flux, $\gamma$-ray pulsar period, radio pulsar period, $\gamma$-ray pulsar $\dot{P}$, radio pulsar $\dot{P}$, $\gamma$-ray pulsar galactic latitude, radio pulsar galactic latitude, $\gamma$-ray pulsar galactic longitude and radio pulsar galactic longitude.
@@
Next, we use statistical tests to examine how well the predicted parameter distributions match the observed distributions for each physical quantity. Given that the simulation and catalog have different-sized samples on each quantity, and the quantity of each pulsar should be independent of each other, we use Mann-Whitney U test (MWU test; Mann & Whitney, 1947) to make the comparison. The MWU test is chosen because it is a nonparametric test independent of underlying distributions. The null hypothesis is that both the predicted and observed samples are drawn from the same underlying distribution. Given the model and simulation have different size, and there is no certain theory on the shape of the distribution of each parameter, it is important to use a statistical test that is independent of the distribution itself. We record the $p$-value of the test for a given physical quantity between the comparison of simulation and catalog. If $p<0.05$, it means we reject the null hypothesis that the two compared samples are drawn from the same distribution. When $p>0.05$, we cannot reject the null hypothesis, which means we consider the two distributions to be the same.
@@
Table 5: P-value of Mann-Whitney U Test, comparing the physical quantities obtained from simulations and catalog. Quantities listed in the second to the third last columns are, respectively, radio flux, $\gamma$-ray flux, period of radio pulsars, period of $\gamma$-ray pulsars, $\dot{P}$ of radio pulsars, $\dot{P}$ of $\gamma$-ray pulsars, $\dot{E}$ of radio pulsars, $\dot{E}$ of $\gamma$-ray pulsars, Galactic latitude of radio pulsars, Galactic latitude of $\gamma$-ray pulsars, Galactic longitude of radio pulsars, and Galactic longitude of $\gamma$-ray pulsars. The last column is the number of the quantities that have p-values $>0.05$ for each model. For a given model, the p-values for each parameter is determined using the random realisation of the model that is shown in Fig. 2
Model $\dot{E}$(G) $\dot{E}$(R) Flux (G) Flux(R) Period (G) Period(R) $\dot{P}$(G) $\dot{P}$(R) b(G) b(R) l(G) l(R) NO
INIT 0.059 0.09 0.0 0.0 0.001 0.0 0.999 0.0 0.767 0.487 0.144 0.052 7
D1 0.58 0.526 0.0 0.0 0.0 0.0 0.443 0.0 0.085 0.541 0.048 0.011 5
D2 0.0 0.0 0.0 0.0 0.001 0.0 0.05 0.0 0.548 0.728 0.426 0.017 4
D3 0.113 0.528 0.0 0.0 0.003 0.0 0.644 0.0 0.461 0.085 0.112 0.047 6
B1 0.019 0.094 0.0 0.0 0.003 0.4 0.266 0.0 0.117 0.579 0.041 0.006 5
B2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.845 0.572 0.058 0.064 4
P1 0.524 0.0 0.0 0.0 0.161 0.0 0.497 0.0 0.819 0.515 0.988 0.003 6
P2 0.758 0.363 0.0 0.0 0.0 0.0 0.241 0.0 0.242 0.893 0.388 0.073 7
P3 0.097 0.0 0.0 0.0 0.029 0.0 0.646 0.0 0.04 0.405 0.003 0.124 4
PN1 0.379 0.003 0.0 0.0 0.66 0.0 0.181 0.0 0.405 0.074 0.267 0.013 6
PN2 0.312 0.09 0.0 0.0 0.006 0.0 0.295 0.0 0.315 0.589 0.019 0.008 5
PN3 0.193 0.858 0.0 0.0 0.008 0.0 0.031 0.0 0.159 0.265 0.624 0.031 5
BLN1 0.191 0.022 0.0 0.0 0.007 0.005 0.048 0.0 0.868 0.047 0.53 0.084 4
BLN2 0.038 0.042 0.0 0.0 0.0 0.0 0.0 0.0 0.268 0.035 0.014 0.005 1
BLN3 0.006 0.0 0.0 0.0 0.908 0.0 0.0 0.0 0.292 0.876 0.121 0.107 5
BLN4 0.84 0.029 0.0 0.0 0.0 0.0 0.0 0.0 0.535 0.159 0.102 0.019 4
PN1-B1 0.618 0.004 0.0 0.0 0.687 0.266 0.065 0.0 0.296 0.836 0.189 0.007 7
BLN1-PN1 0.626 0.002 0.0 0.0 0.54 0.001 0.017 0.0 0.476 0.109 0.151 0.034 5
BLN1-PN5 0.0 0.0 0.0 0.0 0.0 0.0 0.759 0.0 0.654 0.178 0.283 0.001 4
BLN4-PN1 0.0 0.002 0.0 0.0 0.102 0.0 0.0 0.0 0.224 0.086 0.031 0.02 3
In Table. 5, the p-values of the MWU test for a selection of physical quantities of the pulsars compared between each model and the catalog are shown. For each model, the p-values shown in Table. 5 are determined from the random realisation of the model shown in Fig. 2. We note a few of the prominent features in this table below:
  1. 1.
    None of the given models can reproduce the distribution of $\dot{P}$ of radio pulsars. While the CDFs of radio pulsar $\dot{P}$ look similar between the catalog and some of the models as shown in Fig. 2, the MWU p-values are always zero.
  2. 2.
    None of the given models can reproduce the distributions of radio flux. This is surprising, as the distribution of radio luminosities that we assume is based on radio observations (Szary et al., 2014). This is likely due to the fact that the models we explored in this study do not cover a larger or fine enough grid within the parameter space, which is unlike some of the commonly used Bayesian analysis methods or machine learning (Graber et al., 2023; Ronchi et al., 2021) that can explore a larger range of the parameter space with the ability to comb through different parts finely. It is likely that a slightly different choice of radio luminosity distribution would yield a better fit to the flux distribution since this is not something that we varied in our models. Additionally, the catalog is made up of pulsars from various pulsar surveys, which each individually would pose different selection effects on the radio pulsar population. In our work, we only chose to implement selection effects based on the Parkes multi-beam pulsar survey (cf. Section 2.7.2), and this should also have an impact on the resulting radio flux distribution.
  3. 3.
    Similarly for $\gamma$-ray flux, all of the models report a MWU test p-value of 0. When examining the CDFs of the $\gamma$-ray flux (e.g. in Fig. 3), we find that the simulated $\gamma$-ray flux of the pulsars peaks at a lower value compared to those of the catalog. This effect is to some level expected as the changes from the sensitivity level mean that our simulation might be sensitive to $\gamma$-ray pulsars with lower flux. Given that the difference in the peak value is roughly a factor of 2, we think this is an acceptable result as it is important to keep the selection of $\gamma$-ray sensitivity as described above to reflect the observations. Additionally, a recent conference abstract (Tabassum & Lorimer, 2024) suggested that the 3D fundamental plane might not work well for $\gamma$-ray pulsars. As described in § 2.8.1, we use 3D fundamental plane in this work to model the $\gamma$-ray luminosity of pulsars. This might also cause problem in matching model to the catalogs.
  4. 4.
    Only one model, which is the PN1-B1 model, is able to reproduce the period distribution of radio pulsars that resemble that of the catalog.
  5. 5.
    In general, our models reproduce $\gamma$-ray pulsar populations that have a better resemblance to the catalog, compared to the modelled radio pulsar populations. This means that more physical quantities for $\gamma$-ray pulsars have p-values $>0.05$ compared to radio pulsars.
  6. 6.
    The Galactic coordinates of the pulsar populations seem to be better fitted compared to the other physical quantities. This seems to confirm that the Galactic model, kick distribution and other related models used in this work is good enough. However, the large distance uncertainties of the pulsars in the catalog might be a restraint on this modelling.
Since none of the COMPAS models produce results that completely agree with observations, we will not attempt to choose a best-fit model. Here we examine and discuss a few models that perform better than the rest by checking the number of quantities that have p-value $>0.05$. The PN1-B1, P2, and INIT models have 7 quantities with large p-values; and the P1, PN1 and D3 models have 6 quantities with large p-values. We examine these models in detail below, combining the results from Table. 3 in the discussion. In addition, we also pay attention to the B1 and BLN1-PN4 models, as discussed below. From Table 4, it is obvious that some models are badly matched to the observations, including the D2, B2, P1, P3, PN1, BLN2, BLN3, BLN4, BLN1-PN1, and BLN4-PN4. The bad match for model D2 means that short magnetic decay timescale is not favoured. The bad match for models B2, BLN2, BLN3, and BLN4, means that low birth magnetic field is not favored. The bad match for models P1, P3 and PN1 means that large birth rotational periods are not favoured. On the other hand, simply by examining this table, it appears that the model’s INIT, D1, D3, B1, P2, BLN1, and PN1-B1 match the catalogs. Combining the results in Tables. 4 and  5, it appears that models INIT, D3, PN1-B1 can best describe the observed properties of canonical pulsar populations. Examining the CDFs on the physical parameters of all 10 realisations of the INIT, D3, and PN1-B1 models, it appears that visually, the realisations of the PN1-B1 model best match to the distributions of the physical parameters out of the three models.

### v2

@@ 3.2 Results: Radio Pulsars
Figure 1: CDFs of physical quantities for radio pulsars produced by the COMPAS models described in Table 1 (coloured), compared to the data from the catalogues (black). CDFs from a given model are drawn from a randomised sampling from 10 realisations. From top left to bottom right, the panels show the distributions of $\dot{E}$, radio flux, period, $\dot{P}$, Galactic latitude, and Galactic longitude.
@@
We use statistical tests to examine how well the predicted parameter distributions match the observed distributions for each physical parameter of the pulsar populations. These physical parameters include: radio flux, $\gamma$-ray flux, spin period $P$, spin-down rate $\dot{P}$ , spin-down luminosity $\dot{E}$ , Galactic longitude (l) and latitude (b). Given that the observed data and simulations have different-sized samples, and the physical parameters for each pulsar are independent measurements that is not influenced by measuring a different pulsar, we use Mann-Whitney U test (MWUtest, MWU test;) to make the comparison. The MWU test is chosen because it is a nonparametric test independent of underlying distributions. The null hypothesis is that both the predicted and observed samples are drawn from the same underlying distribution. Given the model and simulation have different size, and there is no certain theory on the shape of the distribution of each parameter, it is important to use a statistical test that is independent of the distribution itself. We record the $p$-value of the test for a given physical quantity between the comparison of simulation and catalogue. If $p<0.05$, it means we reject the null hypothesis that the two compared samples are drawn from the same distribution. When $p>0.05$, we cannot reject the null hypothesis, which means we consider the two distributions to be the same.
@@
Table 3: Columns 2 - 7: P-values of Mann-Whitney U Test, comparing the distribution of physical quantities obtained from simulations and catalogue for radio pulsars. Quantities listed in the second to the second last columns are, respectively, radio flux, period, $\dot{P}$, $\dot{E}$, Galactic latitude, and Galactic longitude. The last column: the averaged number of the physical quantities with p-values >0.05>0.05 out of all 10 realisations of each model. For a given model, the p-value for each parameter is determined using the random realisation of the model that is shown in Figure 1.
Model Flux(R) Period(R) $\dot{P}$(R) $\dot{E}$(R) b(R) l(R) NO
INIT 0.0 0.13 0.0 0.0 0.38 0.09 2.4
D1 0.0 0.25 0.0 0.0 0.5 0.52 2.4
D2 0.0 0.0 0.0 0.38 0.35 0.14 2.5
D3 0.0 0.01 0.0 0.0 0.26 0.42 2.1
D4 0.24 0.0 0.0 0.0 0.14 0.01 1.7
D5 0.0 0.0 0.0 0.0 0.35 0.14 2.0
B1 0.0 0.0 0.06 0.0 0.71 0.58 2.3
B2 0.0 0.0 0.0 0.0 0.29 0.18 2.0
P1 0.0 0.0 0.15 0.0 0.4 0.5 2.9
P2 0.0 0.04 0.0 0.0 0.14 0.06 1.9
P3 0.0 0.0 0.57 0.0 0.74 0.48 2.9
PN1 0.0 0.0 0.0 0.0 0.48 0.07 1.6
PN2 0.0 0.09 0.0 0.0 0.54 0.6 2.6
PN3 0.0 0.0 0.0 0.0 0.09 0.34 2.0
PLN1 0.0 0.0 0.0 0.0 0.23 0.13 2.0
BLN1 0.0 0.0 0.0 0.0 0.84 0.48 1.6
BLN2 0.0 0.0 0.0 0.0 0.79 0.19 1.9
BLN3 0.0 0.0 0.0 0.0 0.88 0.06 1.9
BLN4 0.0 0.56 0.0 0.0 0.47 0.21 2.9
P2-D1 0.0 0.0 0.0 0.0 0.75 0.02 1.6
P2-D2 0.0 0.0 0.0 0.01 0.74 0.14 2.6
P2-D3 0.0 0.0 0.0 0.0 0.24 0.0 1.8
P2-D4 0.46 0.0 0.0 0.0 0.8 0.72 2.9
P2-D5 0.0 0.01 0.0 0.0 0.39 0.88 2.0
PN1-B1 0.0 0.0 0.23 0.0 0.47 0.05 2.3
PN1-B1-D1 0.0 0.0 0.0 0.0 0.57 0.36 2.0
PN1-B1-D2 0.0 0.41 0.0 0.0 0.25 0.57 3.0
PN1-B1-D3 0.0 0.0 0.16 0.0 0.07 0.24 2.7
PN1-B1-D4 0.2 0.0 0.0 0.02 0.54 0.62 3.0
PN1-B1-D5 0.0 0.0 0.11 0.0 0.4 0.29 2.7
PN5-B1-D1 0.0 0.0 0.07 0.0 0.18 0.84 1.9
PN5-B1-D2 0.0 0.0 0.01 0.0 0.25 0.56 2.8
PN5-B1-D3 0.0 0.0 0.92 0.0 0.83 0.97 2.8
PN5-B1-D4 0.23 0.14 0.04 0.07 0.96 0.39 3.9
PN5-B1-D5 0.0 0.0 0.27 0.0 0.35 0.35 3.0
BLN1-PN1 0.0 0.0 0.0 0.0 0.86 0.91 1.9
BLN4-PN1 0.0 0.34 0.0 0.0 0.44 0.97 3.0
BLN1-PN5 0.0 0.0 0.0 0.0 0.23 0.92 1.9
BLN5-PLN2-D4 0.21 0.0 0.0 0.74 0.22 1.0 3.4
BLN6-PLN3-D4 0.07 0.0 0.0 0.11 0.68 0.07 3.6
To account for the stochastic variation due to the limited sample size from the models, the observation part of post-processing is repeated 10 times. In Table 3, the $p$-values of the MWU test are shown for a selection of the physical quantities of the pulsars compared between each model and the catalogue. The $p$-values listed in Table 3 are based on the random realisation of the model shown in Figure 1. To further quantify model performance, we introduce the metric NO, defined as the average number of physical quantities (out of six considered) for which the MWU test $p$-value exceeds 0.05 across the 10 realisations. A higher NO value indicates better agreement with the observed distributions, with a maximum possible score of 6. Notable trends and features from Table 3 are discussed below.
The spatial distributions of the simulated radio pulsars closely match those in the catalogue. This is evident as the $p$-values for $b$ across all models, and for $l$ in all but the D4 and P2-D3 models, are larger than 0.05, meaning the simulation results align well with the catalogue distributions shown in Figure 1. This is also an indication that the Galactic models are well-suited for representing radio pulsars.
Only six models successfully reproduce the distributions of radio flux: D4, P2-D4, PN1-B1-D4, PN5-B1-D4, BLN5-PLN2-D4, and BLN6-PLN3-D4. Seven models capture the spin period distribution, nine for the $\dot{P}$ distribution, and four for the $\dot{E}$ distribution. Further examination of the CDF plots for these properties indicates that models B1, BLN1, BLN1-PN1, PN1, D1, D5, P2, PN1-B1, PN1-B1-D2, PN3, PN5-B1-D4, BLN5-PLN2-D4, and BLN6-PLN3-D4 visually exhibit a good fit.
Based solely on the MWU test p-values, model PN5-B1-D4 best matches the observed radio population. It achieves an average NO score of 3.9, meaning that on average, 3.9 out of 6 physical quantities have p-values exceeding 0.05 across the 10 realisations. The next best-performing models are BLN6-PLN3-D4 with NO score of 3.6, BLN5-PLN2-D4 with NO score of 3.4, then BLN4-PN1 and PN1-B1-D4 each with NO score of 3.0. Close behind are models PN5-B1-D5, PN1-B1-D2, P2-D4, P3, P1, and BLN4, each with an average of 2.9 quantities showing p-values above 0.05. Interestingly, despite good visual comparisons, models PN1-B1-D4, BLN4-PN1, P2-D4, P3, P1 and BLN4 do not exhibit a high number of physical quantities p-values exceeding 0.05. All the models mentioned earlier in this paragraph present a mixture of long (PN5 series models and P3) and short (PN1 series models) initial spin periods; large (INIT, $\tau_{d}=500$ Myr series models) and small (D2/D4 series models) magnetic decay timescale. All the models mentioned earlier in this paragraph favour larger initial magnetic field strengths.
3.3 Results: $\gamma$-ray Pulsars
The results of modelling $\gamma$-ray pulsars, and the comparisons to the catalogue are presented in this section. The catalogue $\gamma$-ray pulsars are chosen from 4FGL-DR4, the 14-Year Version of the 4th Fermi-LAT Source catalogue [^6] https://fermi.gsfc.nasa.gov/ssc/data/access/lat/14yr_catalog/ 4FGL-DR4. Only Galactic $\gamma$-ray pulsars not associated with any globular clusters that have $P>0.03$ s and $\dot{P}>10^{17}$ s/s are chosen for model comparison. To ensure consistency, both the catalogue and simulated $\gamma$-ray pulsars are selected from the regions covered by the Parkes Multibeam survey. The simulated $\gamma$-ray pulsars are compared to the catalogue $\gamma$-ray pulsars using the same approach applied to radio pulsars, as described in § 3.2. For the $\gamma$-ray pulsars, the same random realisations of each model, as used in § 3.2, are selected to plot the CDFs in Figure 2 and to list the MWU test p-values in Table 4.